Your company currently has $ 1 comma 000 ​par, 5.25 % coupon bonds with 10 years to maturity and a price of $ 1 comma 071. If you want to issue new​ 10-year coupon bonds at​ par, what coupon rate do you need to​ set? Assume that for both​ bonds, the next coupon payment is due in exactly six months. You need to set a coupon rate of nothing​%. ​(Round to two decimal​ places.)
Answer #1

Since the next coupon payment is in 6 months, its a semi annual bond

Coupon = (0.0525 * 1000) / 2 = 26.25

Number of periods = 10 * 2 = 20

Yield to maturity = 4.37%

Keys to use in a financial calculator: 2nd I/Y 2, FV 1000, PMT 26.25, N 20, PV -1,071, CPT I/Y

If the company wants to issue new bonds at par, coupon ate should be equal to yield to maturity

Therefore, company should set coupon rate as 4.37%